OpenStudy (anonymous):

The Earth’s path around the Sun is elliptical, and the Sun is situated at one of the foci. If the Earth’s orbit can be represented by x^2/25 + y^2/16 = 1, where x and y are measured in 100 million miles, what is the distance from the center of Earth’s elliptical path to the sun? A.)100 million miles B.)200 million miles C.)300 million miles D.)400 million miles

OpenStudy (anonymous):

theyre looking for the distance from the center of the given ellipse to one of your foci... since c^2 = a^2 - b^2 where a=5, b=4 can you find c and answer the question?

OpenStudy (anonymous):

but that gave me 1,681

OpenStudy (anonymous):

c^2 = a^2 - b^2 c^2 = 25 - 16 c^2 = 9 c = ???

OpenStudy (anonymous):

Ohh I added, that's where I went wrong, but c=81

OpenStudy (anonymous):

no trye agen

OpenStudy (anonymous):

c = 3 3*100 million miles = 300 million miles?

OpenStudy (anonymous):

yes... you got it... :)